OpenStudy (anonymous):

Rectangular garden 60 ft x 80 ft surrounded by a walk of uniform width. The area of the walk is 7,200 ft^2. Find the width, x, of the walk.

OpenStudy (anonymous):

since the garden is surrounded on all sides by a walk of uniform width x then the outer length becomes x+80+x = 80+2x and outer width becomes x+60+x = 60+2x area of the walk = outer area - inner area thus we have 7200 = (80+2x)*(60+2x) - 60*80 7200 = 80*60 +2x*80 + 2x *60 +4x^x -60*80 i hope u can do further...